Shift summary for owl 11/27/99

Jason Hamm (jhamm@physics.arizona.edu)
Sat, 27 Nov 1999 06:00:47 -0700 (MST)


Shift Summary for owl shift, 11/27/99

On Shift: R. Pordes, J. Hamm

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 8.25
Hours of E799 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): .25

Accounting of lost time: Beam aborts, fresh start, setting up for muon
run.
========================================================================
Runs ended this shift:

15064 - BEAM_799 (26M events)
15065 - MUON (~100K events)

Tapes brought to FCC this shift: UP[A,B,C] 197-199, UPS066
========================================================================
Accesses during shift: None
========================================================================
Overview of shift: 799 data taking
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15064 01:00 Triggers stop due to DDD 1,2,4 vetos. DAMP
monitoring had frozen earlier in the run, we
assume the tapes are full. Stop process hangs, so
we fresh start. Also do a SDAQ fresh start, as
requested by previous shift.

01:20 Start muon run 15065.

01:25 Stop run 15065 after ~100K events. Flip NM4AN
polarity to positive.

01:37 Start BEAM_799 run 15066.

15066 55 02:47 Missing FFFF word in stream 5 - kfrend.

03:25 Onsplit requests UPF050 on ktev1.

05:00 Onsplit requests UPH035 on ktev4.

06:00 Onsplit requests UPK026 on ktev2.

06:30 Onsplit xterminal crashes. Reboot, restart
xtape and onsplitw.
SDAQ error message claims that scaler_check has
died. Restart all SDAQ processes.

15066 240 06:55 Missing FFFF word in stream 5 - kfrend.

MONITORING ERRORS THAT SHOULD BE FOLLOWED UP ON:

- CsI channel 166 has comparator problems in both 799 runs tonight.
- DC2 plane 4 has high multiplicity (5.1)
- SA4 still has low ADC count distribution. SL has traced this to the SA4
pulse being slightly out of time with the FERA ADC gate.
- HCC bit 2168 is inefficient below 2 GeV, probably due to the poor
dynode signal entering ET.

========================================================================
Follow up Notes on Previous Shift

Monitoring gain drift for channels 1476 and 2893...

Start of shift End of shift
-------------- ------------
1476 -5.30% -5.42%
2893 -5.05% -5.00%

========================================================================
Items to Follow up on Next Shift

========================================================================